Back to the NIKA2 Pool page

Daily Reports

Feb. 12th (Tuesday)

Bilal, Alessia, Denis, Marco, Stefano, Nico arrived.

[Alessia took the 1st shift from 4pm to midnight.]

Start of the observations just after the maintenance, everything went smoothly. At the beginning of the observations the FWHMs were a bit large and we decide to perform observations on primary and secondary calibrators for calibration purpose. We then followed GRB190114C (D02-18) for 2 hours. After the sunset the beams went back to nominal values and we started the observations of the LPSZ.

[Night Shift, Marco and Stefano]

The sunset today gave us good hope:


[This partial illumination of the primary mirror by the setting sun, is surely causing the distortions of the beams often seen with NIKA2 in the afternoon and evening hours. (CK)]
[Details about the foto, for information: the foto was taken at 18:46 local time. At that time the telescope was observing at Az~169 and El~25 (see Logbook). The sun was at Az~253 and El~0. Thus, using the cosine law, the distance from the sun turns out to be 84.56 deg. The sun affects surely the beam during daytime, but for once this probably was not the case at that time. I've seen much worse positions during other pools. (SB)]

We take over at midnight sharp, during the 5th repetition of PSZ2G133 for program 199-16. We repeat it till the 9th repetition (included), taking care of performing the appropriate pointing and PSF checks.

After the usual skydip, pointing, focus, we move to program 192-16 and observe 5 repetitions of the COSMOS scan/cross-scan, with further pointing/PSF checks in between.

We then move further to 204-18. After re-focusing we perform 2 repetitions of NEP_L and one repetition on NEP5, with the usual pointing/PSF checks in between.

Unfortunately we lose 10 minutes at the beginning of 204-18 observations for technical reasons (see Tapas).

Tau(225) is stable at ~0.3 until 3 a.m., then spiky for roughly one hour, and finally stable again at ~0.25 later on.

Probably the creation of IMBFITS stopped on 20190213s5: at 5 a.m. UT, the PIIC monitor is inactive since 272 minutes. This scan was also refused by the IDL RTA and Nicolas reports that it is not properly written.

The new NK_RTA_2 script works very well.

Feb. 13th (Wednesday)

[8am: Denis and Nico are taking their shift.]

- let the loop on NEP finish

- skydip

- MWC349, CRL2688, NGC7027

- Starting GASTON 8h30UT

- done with Gaston (11h30 UT). We wanted to go for a skydip, beammap and calibrators but the beams are 1mm are 14 arcsec, so we move to a time filler program that is ok with the 18 arcsec beam at 2mm : 176-18 / Ucam (DRser is now to low). We could complete this project this afternoon if everything goes fine. We'll keep monitoring the beam regularly.

- 14h UT: beams are too large, we switch to EMIR.

[Afternoon: Alessia shift.]

- After setting up observations for EMIR we have realized that there was a mistake with the visibility range, all the targets for both projects are visible only during the night. We have informed the EMIR pool manager WK and we came back to NIKA2 observations. Beams came back to normal so we finished the observations on the UCAM target for the project 176-18 and we have performed one scan of NGC7538_CS (159-18) before going to Uranus, check the focus and do a beammap [scan 20190213s181]. Observation on the project 123-18 and COSMOS for the 192-16 project (Deep field).

[Night Shift: Marco and Stefano]

The sunset today:


At midnight we start our shift at the telescope. Now we have the COSMOS in our hands; thanks Alessia for handing it over (“lascio il cosmo nelle vostre mani” cit.).

We complete the 4th repetition of the day, then pointing, focus, skydip, pointing and continue with 2 other repetitions of COSMOS (project 192-16).

We switch to project 199-16 and we perform 7 repetitions of source PSZ2G133. This completes the 16 total rep. requested.

Afterwards we perform 3 repetitions of NEP_L for 204-18, with the usual pointing/PSF checks. So far the total on this field is 5 rep.

Pointing and focus were very stable during the whole shift.

Tau was stable at ~0.15 the whole time, except few small spikes. Around 1 a.m. UT the tau fitting process stopped and we figured it out only after some time, sorry. The process was then restarted by Frederic, our operator.

The nk_rta_2 process today stopped few times.

Scans 20190214s4 and s45 seem to be bugged. The makeimbfits stopped and it was restarted following Albrecht’s instructions.

Feb. 14th (Thursday)

[Morning shift: Denis and Nico]

We did a skydip and secondary calibrators before moving to project 111-18 on Saturn's satellites

[Afternoon: Alessia takes over.]

14h LT beams start to degrade and we switch to EMIR observations until 19h00. Second good beammap on Uranus scan # 20190214s214. Start of the project 159-18 but delayed because of a problem with VAC5. ~1h on 159-18 after rebooting VAC5, one scan with bloc manquant. Move to project 123-18 for ~1h before going to COSMOS.

[Night Shift: Marco and Stefano]

While waiting for our shift, here comes today's "Picture from Veleta":


We start our shift shortly after 23:00 local time, during the first repetition of the day on COSMOS (192-16). We continue for 6 repetitions in total.

Around 00:40 the alarm for the cryostat trap P101 rings. The detectors temperatures is still stable. Pressure is increasing

We switch to project 199-16 and integrate for almost 4h on source PSZ2G081 (i.e. 9 repetitions).

We finally start observing the source DRSer of project 176-18, right before the shift changes.

Pointing and focus are repeated regularly. Today they're not as stable as yesterday and we do them more often

Tau is ~0.2 the whole night, with some slow trend at increasing.

The new nk_rta_3 (IDL) does not go through midnight

makeimbfits crashed a couple of times and was re-activated.

We instruct the observers of the morning shift on how to verify the percentage of recovered flux of Calibrators using the PIIC Monitor. This will allow them to check against flux (calib) "jumps" on the fly.

Feb. 15th (Friday)

[Morning shift: Denis and Nico]

The morning shift went pretty well, nothing special to report.

In the early afternoon we noticed that the beams were degrading like the past days, so we switched to EMIR.

Around 2pm - 2.30pm, Santiago fixed last night's issue with the trap.

[Afternoon Shift: Alessia]

EMIR observations until ~19h00 LT. Then observations on Uranus for calibration and project 159-18. Stopped at 20h15 LT because of high wind.

[Night Shift: Marco and Stefano]

The horizon at 19:00...


...but the wind cleaned the sky (even too much).

We start the shift at 23:00 (local) with the telescope parked in safe position because of the strong wind. After half an hour we are able to start observations. Start the DAQ, tuning (by Alessia), setup and then we observe COSMOS (project 192-16) and PSZ2G081 (199-16). Skydip, pointings and focuses in between. Tau(225)~0.15 and lower. We stop at 3:49 a.m. UT because of strong wind again.

The nk_rta_3 script makes it through midnight :-) The makeimbfits procedure kept stopping every few scans, before midnight UT. Afterwards all went smooth.

Feb. 16th (Saturday)

[Stopped because of the wind and working in the meanwhile :)]

First preliminary images of the wide-field in the high-mass star forming galactic region NGC7538 (project 159-18).


[Night Shift: Marco and Stefano]



The telescope is parked. At 23:35 UT it seems that the wind lets us start. We setup all softwares (pako, IDL RTA, make-imbfits, PIIC monitor, etc); we tune the KIDs; pointing, focus, and finally start observing COSMOS (192-16). Unfortunately we have to abort already at the first scan, because the wind is too strong again. It's 00:11:18 UT.

We start observing again at ~4 UT, same procedure as before and then project 199-16, object PSZ2G081 for ~1h. Skydip and finally we switch to 128-18 and start the observation of source L134.

Feb. 17th (Sunday)

[Morning shift: Denis and Nico]

7am UT: we're going for Gaston (L24p5 etc...). Note: OphL1688 is very low on the horizon and clouds are coming up. 9am UT: skydip, calibrators NGC7538 (159-18): two scans before leaving the antenna for a Xpol project 11am UT

[Evening shift: Alessia]

18h00 UTC. Start smoothly observations after XPOL session. Stable weather conditions. Tau below 0.3. Calib map on Uranus before going to project 159-18. Then project 123-18 and start of COSMOS (192-16).

10pm Local TIME: CRASH with error message 'Elvin'... We followed the procedure designed last time, 1st rebooting mrt-lx3, then killing jobs on mrt-lx1... It was not enough. We then rebooted nika2-a with Alain assisting us by Skype. Eventually we managed to restart everything and to resume observations around 23h LT

[Night Shift: Marco and Stefano]

A timid sun today showed its green rim:


We start observing at ~22:30 UT. We observe COSMOS (proj. 192-16) for roughly 2 hour, followed by ~1h15m on PSZ2G081 (proj. 199-16). Skydip, pointings, focus were performed in between and regularly.

We stop observations at 2:31 UT because of strong wind again.

We start again at 5:45 UT. Tuning, Pointing, Focus, and we start to observe Oph L1688 (122-16).

Tonight the PIIC monitor does not process COSMOS, nor PSZ2G081 data and gives the message "data with problems; will not process". It processes pointing data without any problem. The data do not seem to be corrupted: the IDL RTA does not encounter any problem.

[This is the same situation as for DAQ v2&3, when RTA "did not encounter any problem" neither; RZ]
[Yesterday and before the PIIC monitor did reduce all scans, avoiding only the very long ones and saying "scan XXX seconds, processing would take too long". I don't know what might be different today. SB]

Feb. 18th (Monday)

[Morning shift: Denis and Nico]

Stop observations at 10h30 am because of the wind.

[Afternoon shift: Alessia]

Recover NIKA2 observations at 15h00 UTC. Calib scan on Uranus and secondary calibrators before starting the project 209-18. Today we have a foggy sky and even at 15h00 UTC the beams at 1mm and 2mm look pretty nice. This could be an indication of the strong impact, on the broadening of the beams, of temperature and water vapor produced by the melting of the ice in case of sunny days. [Diffuse light caused by fog or a high overcast makes that the dish is more evenly illuminated. It is the strong partial illumination of the primary that can cause the primary to distort. CK]

18h30 UTC: we stop here, it starts to snow

Obsrervations start again at ~21:30 UT

[Night Shift: Marco and Stefano]

One week of observations at Pico Veleta:


We take over at ~22:30 UT. The night is foggy and cloudy. We observe targets/projects that accept tau up to 0.5, following the indication of the Pool Manager:

- source B10 of project 148-18
- sources JINGLE 73, JINGLE 27, JINGLE 23 of project 184-18

At 3:46 UT it's snowing and we have to park the telescope, interrupting scan 20190219s105. We wait until the end of the shift, with not much success.

Feb. 19th (Tuesday)

Snow has stopped and opacity is stable around 0.4 which is tolerated by 184-18, so here we go. Note that i've put more stringent requirements on the RTA plots (please remember these are "quicklook" plots and that they do not accurately represent what could be a final analysis), so more kids are flagged out on the monitoring plots. It does not mean that there is a new problem compared to before.

DailyReportsNika2Pool120219 (last edited 2019-02-19 08:20:49 by NikaBolometer)